OpenStudy (anonymous):

Jack spends 1 1/4 as long on his homework as jill. last week, jill spent 7 1/2 hours doing her homework. how long did jack spend doing his homework?

OpenStudy (anonymous):

your job is to multiply \[1\tfrac{1}{4}\times 7\tfrac{1}{2}\] make them improper fractions and then multiply straight across to get \[\frac{5}{4}\times \frac{15}{2}=\frac{75}{8}=9\tfrac{3}{8}\]